# Behind the Myth of Exploration in Policy Gradients

### v2

@@ Abstract
Policy-gradient algorithms are effective reinforcement learning methods for solving control problems. To compute near-optimal policies, it is essential in practice to include exploration terms in the learning objective. Although the effectiveness of these terms is usually justified by an intrinsic need to explore environments, we propose a novel analysis with the lens of numerical optimization. Two criteria are introduced on the learning objective and two others on its stochastic gradient estimates, and are afterwards used to discuss the quality of the policy after optimization. The analysis sheds the light on two separate effects of exploration techniques. First, they make it possible to smooth the learning objective and to eliminate local optima while preserving the global maximum. Second, they modify the gradient estimates, increasing the probability that the stochastic parameter updates eventually provide an optimal policy. These effects are illustrated empirically on exploration strategies based on entropy bonuses, highlighting their limitations and opening avenues for future works in the design and analysis of such strategies.
Reinforcement Learning, Policy Gradients, Exploration
@@ 1 Introduction
Many practical problems require making sequential decisions in environments. Reinforcement learning (RL) is a framework for solving such decision-making problems that has been successful on complex tasks, including playing games (Mnih et al., 2015; Silver et al., 2017), operating power systems (Aittahar et al., 2024), controlling robots (Kalashnikov et al., 2018), or interacting with electricity markets (Boukas et al., 2021).
@@
Reinforcement learning algorithms interact with an environment to gather information about this environment, which in turn enables to compute and follow an optimal policy. This creates a trade-off between exploration and exploitation. In short, in order to eventually compute a good policy, it is necessary to obtain additional information about the environment by taking actions that are likely not optimal. In algorithms where the trade-off is explicit, exploration is well-understood and has been the subject of many works (Dann et al., 2017; Azar et al., 2017; Neu & Pike-Burke, 2020). In policy-gradient algorithms, one can most often not distinguish exploration from exploitation. Nevertheless, a main theoretical requirement to converge towards globally (or even locally) optimal solutions is that policies remain sufficiently stochastic during the learning procedure (Bhandari & Russo, 2019; Bhatt et al., 2019; Agarwal et al., 2020; Zhang et al., 2021a; Bedi et al., 2022). Interestingly, neither softmax nor Gaussian policies guarantee enough stochasticity for ensuring (fast) convergence (Mei et al., 2020, 2021; Bedi et al., 2022). This requirement of stochasticity in policy-gradient algorithms is often abusively called exploration and understood as the need to infinitely sample all states and actions.
@@
Practitioners have tried to meet the theoretical requirement of sufficient randomness of policies in policy gradient via reward-shaping strategies, whereby a learning objective that promotes or hinders behaviors by providing reward bonuses for some states and actions is optimized as a surrogate to the return of the policy. These bonuses typically promote actions that reduce the uncertainty of the agent about its environment (Pathak et al., 2017; Burda et al., 2018; Zhang et al., 2021c), or that maximize the entropy of states and/or actions (Williams & Peng, 1991; Bellemare et al., 2016; Haarnoja et al., 2019; Hazan et al., 2019; Lee et al., 2019; Islam et al., 2019; Guo et al., 2021; Zhang et al., 2021b). Optimizing a surrogate objective is particularly effective for solving tasks with complex dynamics and reward functions, or with sparse rewards.
@@
The differences between theory and practical implementations of exploration has led to common folklore seeking to provide intuition for the efficiency of policy-gradient methods. This work is part of the research line that studies the maximization of practical surrogate learning objective functions from a mathematical optimization perspective. Close to our work, studies of the learning objective with entropy regularization (an exploration-based reward shaping technique where the entropy of the policy is added in the learning objective) were conducted. It includes the study by Ahmed et al. (2019) concluding that it helps to provide smooth learning objective functions. The same exploration strategy was reinterpreted as a robust optimization method by Husain et al. (2021) and equivalently as a two-player game by Brekelmans et al. (2022). Bolland et al. (2023) furthermore argued that optimizing an entropy regularized objective is equivalent to optimizing the return of another policy with larger variance. More general studies on the learning dynamics have focused on the influence of baselines in policy gradient (Chung et al., 2021), and reward-shaping strategies that do not modify the learning objective, called potential based (Ng et al., 1999; Wiewiora et al., 2003; Harutyunyan et al., 2015; Forbes et al., 2024). All these studies are too restrictive and the literature lacks unified explanations and interpretations about exploration in policy gradients.
@@
Before delving into our contributions, we recall that the convergence of stochastic ascent methods is driven by the objective function and how the ascent directions are estimated. First, the objective function shall be (pseudo) concave to find its global maximum (Bottou, 1998). Second, the convergence rate is influenced by the distribution of the stochastic ascent estimates (Chen & Luss, 2018; Ajalloeian & Stich, 2020). In this paper, we rigorously study policy-gradient methods with exploration-based reward shaping through the lens of these two optimization theory aspects. To that end, we first introduce two new criteria that relate the return of a policy to the learning objective with exploration bonuses, and their respective optima. Second, we introduce two additional criteria on the distribution of the gradient estimates of the learning objective and their likelihood of providing directions in which the learning objective and the return increase. Importantly, these criteria are general to any reward-shaping strategy, and highlight the importance of reward shaping that modify the optimal control behavior, in opposition to the literature on potential-based reward shaping. The influence of some exploration bonuses are illustrated and discussed in the light of these four criteria. In practice, finding good exploration strategies is problem specific and we thus introduce a general framework for the study and interpretation of exploration in policy-gradient methods instead of trying to find the best exploration method for a given task.
@@
The paper is organized as follows. In Section 2, we provide the background about policy gradients and exploration. Section 3 focuses on the effect of exploration on the learning objective while Section 4 is dedicated to the effect on the gradient estimates used in the policy-gradient algorithms [^1] Experimental details and implementations can be found at
https://github.com/adrienBolland/micro-rl-lib.
. Conclusions and future works are discussed in Section 5.
@@ 2 Background
In this section, we introduce Markov decision processes and policy gradients with intrinsic exploration.
@@ 2.1 Markov Decision Processes
We study problems in which an agent makes sequential decisions in a stochastic environment (Sutton & Barto, 2018). The environment is modeled with an infinite-time Markov decision process (MDP) composed of a state space $\mathcal{S}$, an action space $\mathcal{A}$, an initial state distribution with density $p_{0}$, a transition distribution (modeling the dynamics) with conditional density $p$, a bounded reward function $\rho$, and a discount factor $\gamma\in[0,1)$. When an agent interacts with the MDP, first, an initial state $s_{0}\sim p_{0}(\cdot)$ is sampled, then, the agent provides at each time step $t$ an action $a_{t}\in\mathcal{A}$ leading to a new state $s_{t+1}\sim p(\cdot|s_{t},a_{t})$. Such a sequence of states and actions $h_{t}=(s_{0},a_{0},\dots,s_{t-1},a_{t-1},s_{t})\in\mathcal{H}$ is called a history and $\mathcal{H}$ is the set of all histories of any arbitrary length. In addition, after an action $a_{t}$ is executed, a reward $r_{t}=\rho(s_{t},a_{t})\in\mathbb{R}$ is observed.
A policy $\pi\in\Pi=\mathcal{S}\rightarrow\mathcal{P}(\mathcal{A})$ is a mapping from the state space $\mathcal{S}$ to the set of probability measures on the action space $\mathcal{P}(\mathcal{A})$, where $\pi(a|s)$ is the associated conditional probability density of action $a$ in state $s$. The function $J:\Pi\rightarrow\mathbb{R}$ is defined as the function mapping any policy $\pi$ to the expected discounted sum of rewards gathered by an agent interacting in the MDP by sampling actions from the policy $\pi$. We call return of the policy $\pi$ the value provided by that function
@@
where $d^{\pi,\gamma}(\cdot)$ is the discounted state-visitation probability (Manne, 1960). In reinforcement learning, we seek to find an optimal policy $\pi^{*}$ maximizing the expected discounted sum of rewards $J$.
@@ 2.2 Policy-Gradient Algorithms
Policy-gradient algorithms (locally) optimize a parameterized policy $\pi_{\theta}$ to find the optimal parameter $\theta^{*}$ for which the return of the policy $J(\pi_{\theta^{*}})$ is maximized. Naively maximizing the return may provide sub-optimal results. This problem is mitigated in practice with exploration strategies, which consist in optimizing a surrogate learning objective $L$ that intrinsically encourages certain behaviors. In this work, we consider reward-shaping strategies where the expected discounted sum of rewards is extended by $K$ additional reward terms $\rho_{i}^{int}$, called intrinsic motivation terms, and optimize the learning objective
@@
$L(\theta)$ $=\frac{1}{1-\gamma}\underset{\begin{subarray}{c}s\sim d^{\pi_{\theta},\gamma}(\cdot)\\ a\sim\pi_{\theta}(\cdot|s)\end{subarray}}{\mathbb{E}}\left[\rho(s,a)+\sum_{i=0}^{K-1}\lambda_{i}\rho_{i}^{int}(s,a)\right]$
$=J(\pi_{\theta})+J^{int}(\pi_{\theta})\;,$ (2)
@@
where $\lambda_{i}$ are non-negative weights for each intrinsic reward and where $J^{int}(\pi_{\theta})$ is the intrinsic return of the policy. The parameter maximizing the learning objective is denoted by $\theta^{\dagger}$, which we distinguish from the optimal policy parameter $\theta^{*}$. Most of the intrinsic motivation terms can be classified in the two following groups.
@@
Uncertainty-based motivations. It is common to provide bonuses for performing actions that reduce the uncertainty of the agent about its environment (Pathak et al., 2017; Burda et al., 2018; Zhang et al., 2021c). The intrinsic motivation terms are then proportional to the prediction errors of a model of the MDP dynamics. The latter model is usually learned.
Entropy-based motivations. It is also common to provide bonuses for visiting states and/or playing actions that are less likely in histories (Haarnoja et al., 2019; Hazan et al., 2019). In this work, we focus on two of these bonuses
$\rho^{s}(s,a)$ $=-\log d^{\pi_{\theta},\gamma}(\phi(s))$ (3)
$\rho^{a}(s,a)$ $=-\log\pi_{\theta}(a|s)\;,$ (4)
@@
where $\phi(s)$ is a feature built from the state $s$. The corresponding intrinsic returns are maximized for policies that visit uniformly every feature, and for policies with uniformly distributed actions in each state, respectively. Note that these rewards require to estimate the distribution over the states and/or actions. Furthermore, they implicitly depend on the policy parameter $\theta$. The second technique is usually referred to as entropy regularization.
@@
In this work, we consider on-policy policy-gradient algorithms, which were among others reviewed by (Duan et al., 2016) and (Andrychowicz et al., 2020). These algorithms optimize differentiable parameterized policies with gradient-based local optimization. They iteratively approximate an ascent direction $\hat{d}$ relying on histories sampled from the policy in the MDP and update the parameters in the ascent direction, or in a combination of the previous ascent directions (Hinton et al., 2012; Kingma & Ba, 2014). For the sake of simplicity and without loss of generality, we consider that the ascent direction $\hat{d}$ is composed of the sum of an estimate of the gradient of the return $\hat{g}\approx\nabla_{\theta}J(\pi_{\theta})$ and an estimate of the gradient of the intrinsic return $\hat{i}\approx\nabla_{\theta}J^{int}(\pi_{\theta})$. In practice, the first is usually unbiased while the second is computed neglecting some partial derivatives of $\theta$ and is thus biased, typically neglecting the influence of the policy on the intrinsic reward.
@@ 3 Study of the Learning Objective
In this section, we study the influence of the exploration terms on the learning objective defined in equation (2). We define two criteria under which the learning objective can be globally optimized by ascent methods, and such that the solution is close to an optimal policy. We then graphically illustrate how exploration modifies the learning objective to remove local extrema.
@@ 3.1 Policy-Gradient Learning Objective
Policy-gradient algorithms using exploration maximize the learning objective function $L$, as defined in equation (2). We introduce two criteria related to this learning objective for studying the performance of the policy-gradient algorithm. First, we say that a learning objective $L$ is $\epsilon$-coherent when its global maximum is in an $\epsilon$-neighborhood of the return of an optimal policy. Second, we call learning objectives that have a unique maximum and no other stationary point pseudoconcave.
@@
where $\theta^{*}\in\text{argmax}_{\theta}J(\pi_{\theta})$ and where $\theta^{\dagger}\in\text{argmax}_{\theta}L(\theta)$.
@@
If the pseudoconcavity criterion is respected, there is a single optimum, and it is thus possible to globally optimize the learning objective function by (stochastic) gradient ascent (Bottou, 2010) [^2] For the sake of keeping discussions simple, the definition of pseudoconcavity is simplified (Mangasarian, 1975), and additional assumptions on the stochastic gradient estimates are neglected.. If the learning objective is furthermore $\epsilon$-coherent, the latter solution is also a near-optimal policy, where $\epsilon$ is the bound on the suboptimality of its return.
Let us finally remind a theorem from Ng et al. (1999).
Consistency Theorem. The learning objective $L$ is $\epsilon$-coherent, with $\epsilon=0$, in any MDP with state space $\mathcal{S}$, action space $\mathcal{A}$ and factor $\gamma$, if, and only if, $J(\theta)=L(\theta)$ for all $\theta$. The intrinsic rewards are furthermore potential based.
This theorem states that there is no MDP-agnostic exploration method that guarantees consistency with $\epsilon$ equal to zero and that modifies the objective function. This type of exploration is only possible with potential-based reward shaping (Ng et al., 1999). In conclusion, if the return is not pseudoconcave, there is a trade-off between the two criteria, which can not be resolved by potential-based exploration.

### v3

@@ Abstract
In order to compute near-optimal policies with policy-gradient algorithms, it is common in practice to include intrinsic exploration terms in the learning objective. Although the effectiveness of these terms is usually justified by an intrinsic need to explore environments, we propose a novel analysis with the lens of numerical optimization. Two criteria are introduced on the learning objective and two others on its stochastic gradient estimates, and are afterwards used to discuss the quality of the policy after optimization. The analysis sheds light on two separate effects of exploration techniques. First, they make it possible to smooth the learning objective and to eliminate local optima while preserving the global maximum. Second, they modify the gradient estimates, increasing the probability that the stochastic parameter updates eventually provide an optimal policy. We empirically illustrate these effects with exploration strategies based on entropy bonuses, identifying limitations and suggesting directions for future work.
@@ 1 Introduction
Reinforcement learning (RL) has been successful in complex decision-making problems, including playing games (Mnih et al.,, 2015; Silver et al.,, 2017), operating power systems (Aittahar et al.,, 2024), controlling robots (Kalashnikov et al.,, 2018), and trading electricity (Boukas et al.,, 2021).
@@
Reinforcement learning algorithms interact with an environment to gather information about this environment, which in turn enables to compute and follow an optimal policy. This interaction creates a trade-off between exploration and exploitation. In short, to eventually compute a good policy, an agent must obtain additional information about the environment by taking actions that are likely not optimal. In algorithms where this trade-off is explicit, exploration is well understood and has been the subject of many works (Dann et al.,, 2017; Azar et al.,, 2017; Neu and Pike-Burke,, 2020). In policy-gradient algorithms, one often can not distinguish exploration from exploitation. Nevertheless, a key theoretical requirement for convergence to globally (or even locally) optimal solutions is that policies remain sufficiently stochastic during the learning procedure (Bhandari and Russo,, 2019; Bhatt et al.,, 2019; Agarwal et al.,, 2020; Zhang et al., 2021a, ; Bedi et al.,, 2022). Interestingly, neither softmax nor Gaussian policies guarantee enough stochasticity to ensure (fast) convergence (Mei et al.,, 2020, 2021; Bedi et al.,, 2022). This requirement for stochasticity in policy gradients is often loosely called exploration and understood as the need to infinitely sample each state-action pair.
@@
The need for sufficient policy randomness has long been known by practitioners, who have attempted to meet this requirement in policy-gradient methods through reward-shaping strategies. In these approaches, a learning objective that promotes or hinders behaviors by assigning reward bonuses to some states and actions is optimized as a surrogate for the policy return. Such bonuses typically encourage actions that reduce uncertainty of the agent about its environment (Pathak et al.,, 2017; Burda et al.,, 2018; Zhang et al., 2021c, ) or that maximize the entropy of states or actions (Williams and Peng,, 1991; Bellemare et al.,, 2016; Haarnoja et al.,, 2019; Hazan et al.,, 2019; Lee et al.,, 2019).
@@
Optimizing surrogate learning objectives that include reward bonuses appears particularly effective for solving challenging tasks, typically those with complex dynamics or sparse rewards. The connection between such objectives and exploration has nevertheless long been limited to intuition, and the lack of theoretical analysis has led to common folklore that seeks to justify their practical efficiency. Research on these objectives has mainly focused on entropy regularization (an exploration-based reward-shaping technique in which the policy entropy is added to the learning objective). This line of work includes the empirical study of Ahmed et al., (2019), which concludes that entropy regularization helps to provide smoother objective functions. The same strategy has been reinterpreted as a robust optimization method by Husain et al., (2021) and, equivalently, as a two-player game by Brekelmans et al., (2022). Bolland et al., (2023) further argue that optimizing an entropy-regularized objective is equivalent to maximizing the return of another policy with larger variance. Finally, convergence rates for policy-gradient algorithms with entropy regularization have recently been derived (Cen et al.,, 2022; Zhang et al., 2021a, ). More general studies of learning dynamics have examined the influence of baselines in policy-gradient methods (Chung et al.,, 2021) and potential-based reward-shaping strategies that do not modify the learning objective (Ng et al.,, 1999; Wiewiora et al.,, 2003; Harutyunyan et al.,, 2015; Forbes et al.,, 2024). All of these studies remain restricted to particular learning objectives and the literature still lacks unified explanations of the improvements intrinsic reward bonuses can provide, the conditions under which they arise, and systematic ways to compare different bonuses, among others. This work aims to take a step toward answering these questions.
@@
Before delving into our contributions, we recall that the convergence of stochastic ascent methods is driven by the objective function and how the ascent directions are estimated. First, the objective function shall be (pseudo) concave to find its global maximum (Bottou,, 1998). Second, the convergence rate is influenced by the distribution of the stochastic ascent estimates (Chen and Luss,, 2018; Ajalloeian and Stich,, 2020). In this paper, we rigorously study policy-gradient methods with exploration-based reward shaping through the lens of these two optimization theory aspects. To that end, we first introduce two new criteria that relate the return of a policy to the learning objective with exploration bonuses, and their respective optima. Second, we introduce two additional criteria on the distribution of the gradient estimates of the learning objective and their likelihood of providing directions in which the learning objective and the return increase. Importantly, these criteria are general to any reward-shaping strategy, and highlight the importance of reward shaping that modify the optimal control behavior, in opposition to the literature on potential-based reward shaping. The influence of some exploration bonuses are illustrated and discussed in the light of these four criteria. In practice, finding good exploration strategies is problem specific and we thus introduce a general framework for the study and interpretation of exploration in policy-gradient methods instead of trying to find the best exploration method for a given task.
@@
The paper is organized as follows. In Section 2, we provide the background about policy gradients and exploration. Section 3 focuses on the effect of exploration on the learning objective while Section 4 is dedicated to the effect on the gradient estimates used in the algorithms. Both sections include illustrations on pathological examples [^1] Implementations can be found at https://github.com/adrienBolland/micro-rl-lib.. Conclusions and future works are discussed in Section 5.
@@ 2 Background
Let us introduce Markov decision processes and policy gradients with intrinsic exploration.
@@ 2.1 Markov Decision Processes
We study problems in which an agent makes sequential decisions in a stochastic environment (Sutton and Barto,, 2018). The environment is modeled as an infinite-time Markov decision process (MDP) composed of a state space $\mathcal{S}$, an action space $\mathcal{A}$, an initial state distribution with density $p_{0}$, a transition distribution (modeling the dynamics) with conditional density $p$, a bounded reward function $\rho$, and a discount factor $\gamma\in[0,1)$. When an agent interacts with the MDP, first an initial state $s_{0}\sim p_{0}(\cdot)$ is sampled, then, at each time step $t$, the agent provides an action $a_{t}\in\mathcal{A}$, leading to a new state $s_{t+1}\sim p(\cdot|s_{t},a_{t})$. Finally, a reward $r_{t}=\rho(s_{t},a_{t})\in\mathbb{R}$ is observed after each action $a_{t}$.
A policy $\pi\in\Pi=\mathcal{S}\rightarrow\mathcal{P}(\mathcal{A})$ is a mapping from the state space $\mathcal{S}$ to the set of probability measures on the action space $\mathcal{P}(\mathcal{A})$, where $\pi(a|s)$ is the associated conditional probability of action $a$ in state $s$. The function $J:\Pi\rightarrow\mathbb{R}$ is defined as a function mapping any policy $\pi$ to the expected discounted sum of rewards gathered by an agent interacting with the MDP by sampling actions from $\pi$. We call return of the policy $\pi$ the value provided by that function
@@
where $d^{\pi,\gamma}(s)=(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}p_{t}^{\pi}(s)$ is the state-visitation measure $\forall s\in\mathcal{S}$, and $p_{t}^{\pi}$ is the state probability after $t$ time steps following the policy $\pi$. In reinforcement learning, an optimal policy $\pi^{*}\in\Pi$ maximizes the expected discounted sum of rewards $J$.
@@ 2.2 Policy-Gradient Algorithms
Policy-gradient algorithms (locally) optimize a parameterized policy $\pi_{\theta}$ to find the parameter $\theta^{*}$ that maximizes the return of the policy. Naively maximizing the return may provide suboptimal results. This problem is mitigated in practice by exploration strategies, which typically consist of optimizing surrogate learning objectives that encourage certain behaviors. In this work, we consider reward-shaping strategies in which the expected discounted sum of rewards is extended by $K$ additional reward terms $\rho_{k}^{int}$, called intrinsic motivation terms, and we optimize the learning objective
@@
$L(\theta)$ $=\frac{1}{1-\gamma}\underset{\begin{subarray}{c}s\sim d^{\pi_{\theta},\gamma}(\cdot)\\ a\sim\pi_{\theta}(\cdot|s)\end{subarray}}{\mathbb{E}}\left[\rho(s,a)+\sum_{k=0}^{K-1}\lambda_{k}\rho_{k}^{int}(s,a)\right]=J(\pi_{\theta})+J^{int}(\pi_{\theta})\;,$ (2)
@@
where $\lambda_{k}$ are non-negative weights for each intrinsic reward, and $J^{int}(\pi_{\theta})$ is the intrinsic return of the policy. The parameter maximizing the learning objective is denoted by $\theta^{\dagger}$, which we distinguish from the optimal policy parameter $\theta^{*}$. Most of the intrinsic motivation terms can be classified into the following two groups.
@@
Uncertainty-based motivations. It is common to provide bonuses for performing actions that reduce the agent uncertainty about an internal environment model (Pathak et al.,, 2017; Burda et al.,, 2018; Zhang et al., 2021c, ). The intrinsic motivation terms are then proportional to the prediction errors of that model, which is typically learned.
Entropy-based motivations. It is also common to motivate agents to visit states and/or select actions that are less likely. It is typically achieved using intrinsic rewards that depend on the policy (Williams and Peng,, 1991; Haarnoja et al.,, 2019) or on the state-visitation measure (Bellemare et al.,, 2016; Hazan et al.,, 2019; Lee et al.,, 2019; Islam et al.,, 2019; Guo et al.,, 2021; Zhang et al., 2021b, )
$\rho^{a}(s,a)$ $=-\log\pi_{\theta}(a|s)$ $J^{a}(\pi_{\theta})\propto\mathbb{E}_{s\sim d^{\pi_{\theta},\gamma}(\cdot)}\left[\mathcal{H}_{a}\left(\pi_{\theta}(a|s)\right)\right]$ (3)
$\rho^{s}(s,a)$ $=-\log d^{\pi_{\theta},\gamma}(z)$ $J^{s}(\pi_{\theta})\propto\mathcal{H}_{z}\left(d^{\pi_{\theta},\gamma}(z)\right)\;,$ (4)
@@
where $z=\phi(s)$ is a feature built from the state $s$, and where $\mathcal{H}_{x}(q(x))$ is the entropy of $q$. The corresponding intrinsic returns, $J^{a}(\pi_{\theta})$ and $J^{s}(\pi_{\theta})$, are maximized for policies with actions uniformly distributed in each state, and for policies that visit every feature uniformly, respectively. Importantly, these rewards require estimating the distribution over the states and/or actions, which implicitly depends on $\theta$. Notice that the first technique is usually referred to as entropy regularization.
@@
In this work, we consider on-policy policy-gradient algorithms, which were, among others, reviewed by Duan et al., (2016) and Andrychowicz et al., (2020). These algorithms optimize differentiable parameterized policies with gradient-based local optimization. They iteratively approximate an ascent direction $\hat{d}$, relying on samples from the policy in the MDP, and update the parameters in that ascent direction, or in a combination of previous ascent directions (Hinton et al.,, 2012; Kingma and Ba,, 2014). For the sake of simplicity and without loss of generality, we consider that the ascent direction $\hat{d}$ is the sum of an estimate of the gradient of the return $\hat{g}\approx\nabla_{\theta}J(\pi_{\theta})$ and an estimate of the gradient of the intrinsic return $\hat{i}\approx\nabla_{\theta}J^{int}(\pi_{\theta})$. In practice, the first may be unbiased, while the second is computed by neglecting some partial derivatives of $\theta$ and is thus biased, typically neglecting the influence of the policy on the intrinsic reward.
@@ 3 Study of the Learning Objective
In this section, we study the influence of the exploration terms on the learning objective defined in equation (2). We define two criteria under which the learning objective can be globally optimized by ascent methods and under which the solution is close to an optimal policy. We then graphically illustrate how exploration modifies the learning objective to remove local extrema.
@@ 3.1 Policy-Gradient Learning Objective
Policy-gradient algorithms using exploration maximize the learning objective function $L$ as defined in equation (2). We introduce two criteria related to this learning objective to study the performance of the policy-gradient algorithm. First, we say that a learning objective $L$ is $\epsilon$-coherent when its global maximum is in an $\epsilon$-neighborhood of the return of an optimal policy. Second, we call learning objectives that have a unique maximum and no other stationary point pseudoconcave.
@@
where $\theta^{*}\in\arg\max_{\theta}J(\pi_{\theta})$ and where $\theta^{\dagger}\in\arg\max_{\theta}L(\theta)$.
@@
If the pseudoconcavity criterion is satisfied, there is a single optimum, and it is thus possible to globally optimize the learning objective function by (stochastic) gradient ascent (Bottou,, 2010) [^2] The literature on policy-gradient algorithms usually relies on the Polyak–Łojasiewicz inequality and gradient dominance (Polyak,, 1963; Lojasiewicz,, 1963; Karimi et al.,, 2016) rather than pseudoconvexity (Mangasarian,, 1975), all of which imply that every stationary point is a global minimum (invexity), from which global convergence of gradient descent methods follows (Ben-Israel and Mond,, 1986; Mishra and Giorgi,, 2008; Bottou,, 2010). Gradient dominance furthermore ensures linear rates of convergence (Karimi et al.,, 2016). Under quasiconvexity, invexity and pseudoconvexity coincide (Mishra and Giorgi,, 2008) and are thus implied by gradient dominance. For the sake of keeping the discussion simple, the definitions of pseudoconvexity and pseudoconcavity are simplified, and additional assumptions on the stochastic gradient estimates are neglected.. If the learning objective is furthermore $\epsilon$-coherent, the latter solution is also a near-optimal policy, where $\epsilon$ is the bound on the suboptimality of its return.
Let us finally recall a theorem of Ng et al., (1999).
Consistency Theorem. The learning objective $L$ is $\epsilon$-coherent, with $\epsilon=0$, in any MDP with state space $\mathcal{S}$, action space $\mathcal{A}$ and discount factor $\gamma$, if, and only if, $J(\pi_{\theta})=L(\theta)$ for all $\theta$. Furthermore, the intrinsic rewards are potential based.
This theorem establishes a trade-off between the coherence and pseudoconcavity criteria. Either an exploration method guarantees coherence with $\epsilon=0$ in any MDP, but then the learning objective equals the return and the pseudoconcavity criterion is violated when the return is not pseudoconcave itself. Otherwise, the pseudoconcavity criterion is satisfied, but then the exploration method is MDP-dependent or coherence is guaranteed at best with $\epsilon>0$.